# Evolving Agents: Interactive Simulation of Dynamic and Diverse Human Personalities

### v1

@@ 2.3. Personality Evolution Architecture
The term ”personality” refers to the sets of predictable behaviors by which individuals are recognized and identified(ghasem-aghaee_cognitive_2007, ). Personality analysis and description are widely studied in the field of personality psychology, a field that proposes well-known personality classification methods such as the Big Five personality traits(john_big_1999, ) and the Myers-Briggs Type Indicator (MBTI)(coe_mbti_1992, ).
Regarding the mechanisms of personality affect behavior, the widely accepted definition of personality refers to cognitive, emotional and character traits that shape behaviors(zuckerman_psychobiology_1991, ). In previous work, researchers believe that personality changes are related to cognitive complexity(ghasem-aghaee_cognitive_2007, ). Another study have proposed cognitive conceptual frameworks including reasoning, retrieval, learning and grounding(sumers_cognitive_2024, ). Further research has summarized the conceptual architecture of agents in more human-like terms of perception, brain, action(xi_rise_2023, ). In research related to agents, Some agents can adjust emotional responses and engage in emotional resonance(hasan_sapien_2023, ). Additionally, some work has explored character portrayal based on large language models(caron_identifying_2022, ). However, current architectural approaches to agent personality are mostly partial simulations, not yet producing simulations of an integrated personality that can influence behavior.
What’s more, in real word human behavior and personality are always evolving over time, which presents a substantial challenge for agent simulation. A realistic approach to representing this changing process is dynamic personality, which considers the variability of personality traits on the basis of changes at the corresponding personality levels. In early research, some works used discontinuous personality update models to achieve the purpose of personality evolution, and some works dynamically simulated personality through the establishment of fuzzy logic agents. These works, trained on traditional models and primarily utilizing a parametric approach to control personality traits, are often constrained by the number of parameters. This limitation results in a lack of realism in the process of agent personality changes. At the same time, different parameter change processes are often directly inferred by the AI model, lacking an interpretive description of the intermediate process, and observers cannot intuitively understand the reasons for character personality changes. As for current work discussing generative model-based agents, they only simulate static personality at a specific time point and have not yet simulated dynamic personality agents that evolve over time.
Figure 2. Evolving Agents architecture is inspired from psychology stimulus-response theory, encompasses two subsystems: Behavior and Personality
@@ 3. AGENT ARCHITECTURE
To enable agents evolution, we first need to understand how human personalities are composed and evolve. The field of psychology already has many rich studies related to this issue(wang_survey_2024, ). Inspired by the stimulus-response theory in behaviorist psychology(shen_hugginggpt_2023, ), we wish for agents to gain stimulation and feedback through interaction with simulation environments, thereby affecting the development of the agent’s personality.
Based on the psychology theory discussed(zuckerman_psychobiology_1991, ), we propose a novel agent architecture - Evolving Agents, which encompasses two subsystems: Behavior and Personality, as shown in figure 2. The Behavior system guides the execution of observable, personalized actions. The Personality system establishes a cognitive mechanism with individual characteristics, capable of continuously updating personality based on feedback based on the external information. During simulation, the Behavior and Personality systems continuously influence each other. The planning and behavior of the Agent will take its own characteristic into comprehensive consideration. Simultaneously, the experiences generated from an agent’s daily actions also impact personality through memory and insight. This process of mutual influence forms a feedback loop pattern, facilitating continuous personality evolution.
In detail, the Behavior system comprises two main modules: Plan and Action. Its execution includes personalized planning for the Agent, social engagements and responses, and behavior execution functionalities. The Personality system contains three main modules: Emotion, Cognition and Character Growth. Its execution involves subjective emotional perception, daily information filtering and memory decay management, comprehensive context-aware idea summarization, and the growing of the Agent’s character. Each module is empowered by large language models like GPT-4 when dealing with its input.
@@ 3.1. Agent Behavior
The Behavior system is responsible for executing planned actions(wang_survey_2024, ) within a simulated environment. Building on the existed architecture of Agent simulation(caron_identifying_2022, )(curry_computer_2023, )(sumers_cognitive_2024, ), which includes planning (Plan)(zhou_least–most_2023, ) and executing plans (Action), we introduce three innovative elements in the Plan Module: Characteristic Plan, Goal-based Mechanism, and Specific Post-Process. In the Action Module, we propose a new method to describe the Agent’s social relationships and enhance the continuity of conversations between Agents.
@@ 3.1.1. Plan Module
Characteristic Plan: We aim to reflect the personality differences of agents in the behaviors of Evolving Agents. Therefore, in the design of the planning module, we consider not only environmental information as inputs but also integrate agents’ character and memory as references for GPT-4 to generate plans. This ensures that each agent can create a schedule that is align with its personality. The specific mechanism by which the Personality system affects the Plan module will be discussed in the following section3.2.
@@
Figure 3. Each agent can create a schedule that aligns with its character. For example, the enthusiastic and open Isabella prefers to leave her own dorm and make social plans compared to the shy Benjamin
@@
Goal-based Mechanism: We aim to construct a mechanism that can easily analyze the variations between agent behavior and personality. Therefore, we referenced relevant perspectives on behavior and motivation from self-determination theory(noauthor_what_nodate, ), considering planning as an expression of behavioral motivation. In our specific design, we defined categories of goals (such as learning, relaxation, exercise, etc.) and required agents to understand and refer to these goals while planning as shown in figure 3. This is to ensure that agents consider the motivations behind each planned behavior, guiding their thought process effectively. Such a goal-based mechanism also aids observers in summarizing user behavior patterns, making it easier to analyze related pattern evolutions. We will discuss this in detail in Chapter 5.
@@
Specific Post-Process:We provide different planning goals with unique post-processes. Taking ”Appointment” as an example, most existing Agent simulation systems use distance-based and probability-based triggering logic for Agent interactions, as is shown in figure 4. We aim to enhance the proactivity of Agents in interacting with others, fostering more ”conscious” and ”prepared” social activities in the simulated world. In the planning phase, Agents can consider their relationships with other Agents and their current state to decide whether to extend social invitations. The invited Agents respond to the invitation, giving reasons for acceptance or rejection. For overlapping appointments, Agents weigh existing social plans against new ones, prioritizing based on personal development benefits. Following this request-response process, both parties adjust their plans based on the outcome.
@@
For example, Sophia wants to meet Isabella at a cafe at 16:00 to discuss, believing that Isabella’s appreciation for creative expression aligns with her own dream of becoming an inspiring and successful writer. Subsequently, Isabella accepts Sophia’s invitation, reasoning that her own preference for in-depth conversations over superficial socializing perfectly complements Sophia’s personality, suggesting that their meeting could lead to deeper and more meaningful discussions. Interestingly, when faced with simultaneous social invitations, even though Isabella initially planned to engage with another friend, Benjamin, during this time, she considers the conversation with Sophia potentially more crucial for her personal development, and thus decides to accept this invitation.Please refer to the image for the detailed process.
@@ 3.1.2. Action Module
Action Mechanism: When the game time reaches the scheduled plan time, the Agent starts executing the specific plan. By considering the current plan and personality information, GPT-4-turbo generates a detailed, objective natural language description of the Agent’s actions regarding the plan. Before executing the plan, the Agent must first check if the environmental conditions meet the plan’s objectives, adopting the item interaction condition logic from Generative Agents. If the target interaction object is occupied, the Agent will replan until conditions are met. Once the environmental object target for executing the plan is secured, the Agent will take over the interaction spot (in social scenarios, reserving adjacent spots for its social peers) and perform the action.
@@
Continously Developing Communication: When two Agents come within a certain distance of each other, there’s a probability of triggering a dialogue, depending on their social tendencies and current state (for example, if an Agent’s plan is social, it will actively initiate a conversation). From a social psychology perspective, such interactions allow Agents to gain insights from others with different personalities, serving as a vital external stimulus for their own personality development. We aim for the depth of conversation between agents to gradually increase, enabling the agents to generate more valuable insights. To achieve this, we have implemented the dialogue memory. Whenever a conversation occurs, the Agents will review the summary of content previously discussed with the current interlocutor. Then, it determines a ”further” new topic based on the previously discussed topics, and discuss around this new topic. This dialogue mechanism enhances the inspiration and viewpoint generation. For example, Isabella and Sophia’s discussions deepened from personal development and societal needs to specific methods for maintaining creativity and mental health under social pressure, the procedure of which is shown in figure 5. This mechanism mirrors real-world communication, where familiarity and discussion depth lead to increasingly inspiring and detailed insights on certain topics.
@@
Figure 5. Agent choose their communication partner and define the topic to talk based on the dialog memory
@@
Abstract Agent Relationship: Evolving Agents does not use parameters or specific categories to describe agents’ relationship with others. We found that doing so would excessively predetermine interactions with certain Agents. We enable large language models to select preferred conversation partners based on agents’ dialogue memory. Simultaneously, we requires agents to provide reasons for their choices. This approach results in more rational and intriguing social choices, enabling agents to openly choose conversation partners and gain diverse perspectives on their dilemmas. For instance, Isabella, who is balancing external expectations with personal aspirations might focus on issues like personal dreams and social responsibility with Sophia, a literary writer interested in social insights, or focus on balancing personal pursuits and intimate relationships with Isla, a romantic filmmaker, as is shown in figure 6. We believe this process can help agents explore more possibilities during their Personality Evolution.
@@ 3.2. Agent Personality
In the Agent Behavior System described above, Agents produce a variety of objective behaviors through interactions with their environment and other Agents. From the perspective of cognitive psychology, these behaviors can be viewed as stimuli. The Agent Personality system is responsible for processing these stimuli, generating feedback, and thus driving the evolution of the Agent’s personality. In the field of psychology, the widely accepted definition of personality encompasses cognitive, emotional and character traits that influence behaviors. We adapt these elements from psychology theory to fit the operational logic of Agent Architecture, ultimately identifying three main modules within the personality system: Emotion, Cognition and Character Growth.
@@ 3.2.1. Character Structure
Five Dimension Based Character Structure: In psychology, character is defined as a set of information that includes habits and tendencies in thought and action(snowden_oxford_nodate, ), which play a crucial role in their behavior. From the perspective of system design, character directly describes an individual’s traits and state.
A structure that can precisely and comprehensively describe characteristics is essential for the effective integration and transmission of personality information to each module within an agent architecture. We constructed a five-dimension Character Structure based on psychological theories, including Basic Information, Current State, Traits(john_big_1999, )(noauthor_introduction_nodate, ), Conflict(noauthor_apa_nodate, ), and Preference(xi_rise_2023, ),as is shown in figure 7. ”Basic Information” encompasses fundamental information about the Agent that does not change, such as gender, age, profession, etc. In the Current State, we detail the agent’s initial life and psychological state. In Personality Traits, we describe the agent based on the Big Five personality dimensions. In conflict, we detail contradictions within the agent’s personality, facilitating personality evolution around these conflicts. To better describe agent’s behavior characteristics, in Behavior Preference we set a six-dimension criteria which is composed of the ultimate goal, long-term goal(derived by ultimate goal), short-term goal(derived by ultimate goal), daily routine(based on agent’s overall personality), hobbies and venue preference. In practice, we initialize specific agents’ personality using GPT-4-turbo based on this Five Dimension Character structure.
@@
Two Versions of Character Structure: To enhance the efficiency and accuracy of the system development process, we employ two versions of the Character Structure within the system: a full version and a summary version. The full version includes complete description, while the summary version condenses and summarizes the full version. In specific modules, we differentiate between utilizing the full version and the summary version based on the variance in information weight. For instance, in the Behavior system’s Plan module discussed earlier, we emphasize the information weight of Behavior Preference from the Character Structure. Therefore, we provide a full description of Behavior Preference to GPT-4-turbo, but only the summary versions for other dimensions. As for other dimensions, we only provide the summary version. This approach helps the model focus more on the alignment between Plan and Behavior Preference, also effectively reducing the generation time for large language models meanwhile.
@@ 3.2.2. Emotion
Timely Experience and Feeling: Human feelings towards the same objective behavior are not constant. Specific subjective experiences and emotional responses are generated in response to external stimuli under particular circumstances. In specific circumstances, outworld’s stimuli affects us to generate emotional behavior and subjective feelings. As a result, our cognitive process such as attention, thinking pattern and decision is influenced, leading us to differentiated perceptions of external information(solso_cognitive_2005, ). Therefore, it’s crucial to build a system that can not only comprehensively consider the current state and perform emotional perception on ongoing behaviors, but also ensure that later cognition process receive information aligned with the character’s current emotional state. At the implementation level, we categorize the Agent’s current emotional state into seven categories. The Emotion module considers the current personality conflicts and other personality information to update emotion based on the behavior being executed. In detail, we request GPT4 to choose the current emotion class with a first-person perspective description of inner feeling. This module allows Agents to generate subjective descriptions that vary and align with their experiences, such as Sophia feeling excited about writing her personal novel due to anticipation of creating her literary work, or feeling anxious and self-doubtful after hearing the technological innovation in the writing process told by Benjamin, as is shown in Figure 8.
@@
Emotion-Based Plan Evaluation: Previous research has noted that humans don’t rigidly follow the plans they set for themselves. In the design of Humanoid Agent architecture, researchers proposed a dual-system structure of emotion and behavior to allow Agents to constantly adapt their plans based on internal feelings.
The Evolving Agents’ Emotion module also supports continuous evaluation and adjustment in the Behavior system’s Plan module. We simplified the adjustment mechanism of Emotion on Plan, using a more systematic decision-making mechanism to trigger plan adjustments. Emotions are categorized on a scale from negative to positive, mapped within a 0-7 interval. Significant emotional shifts during an action (e.g., from excitement to calm, crossing a 3-interval span) indicate substantial emotional fluctuations, as a result of which, agents should reevaluate the forthcoming plans. This straightforward strategy, as is shown in Figure 9, demonstrates believable and dynamic planning capabilities. For example, after a conversation with Sophia, Isabella’s mood shifted from excitement to fear (due to the discussion regarding social responsibilities caused pressure), leading her to adjust her original plan and go to a square for relaxation to calm her emotions.
Figure 9. Agent evaluate and modify existed plans based on the emotional fluctuations

### v3

@@ 2.3. Personality Evolution Architecture
The term ”personality” refers to the sets of predictable behaviors by which individuals are recognized and identified(ghasem-aghaee_cognitive_2007, ; matthews_personality_2003, ; goldberg_structure_1993, ; hampson_personality_2012, ; edwards_measurement_1973, ). Personality analysis and description are widely studied in the field of personality psychology, a field that proposes well-known personality classification methods such as the Big Five personality traits(gosling_very_2003, ) and the Myers-Briggs Type Indicator (MBTI)(coe_mbti_1992, ).
Regarding the mechanisms of personality affect behavior, the widely accepted definition of personality refers to cognitive, emotional, and character traits that shape behaviors(zuckerman_psychobiology_1991, ). In previous work, researchers believe that personality changes are related to cognitive complexity(ghasem-aghaee_cognitive_2007, ). Another study has proposed cognitive conceptual frameworks, including reasoning, retrieval, learning, and grounding(sumers_cognitive_2024, ). Further research has summarized the conceptual architecture of agents in more human-like terms of perception, brain, and action(xi_rise_2023, ). In research on agents, Some agents can adjust emotional responses and engage in emotional resonance(hasan_sapien_2023, ). Additionally, some work has explored character portrayal based on large language models(serapio-garcia_personality_2023, ). However, current architectural approaches to agent personality are mostly partial simulations, not yet producing simulations of an integrated personality that can influence behavior.
Moreover, human behavior and personality always evolve in the real world, which presents a substantial challenge for agent simulation. A realistic approach to this changing process is dynamic personality, which considers the variability of personality traits based on changes at the corresponding personality levels. In early research, some works used discontinuous personality update models to achieve the purpose of personality evolution, and some works dynamically simulated personality through the establishment of fuzzy logic agents. These works, trained on traditional models and primarily utilizing a parametric approach to control personality traits, are often constrained by the number of parameters. This limitation results in a lack of realism in agent personality changes. At the same time, different parameter change processes are often directly inferred by the AI model, lacking an interpretive description of the intermediate process, and observers cannot intuitively understand the reasons for character personality changes. As for current work discussing generative model-based agents, they only simulate static personality at a specific time point and have not yet simulated dynamic personality agents that evolve.
Figure 2. Evolving Agents architecture is inspired by psychology stimulus-response theory, encompasses two subsystems: Behavior and Personality
@@ 3. AGENT ARCHITECTURE
We first need to understand how human personalities are composed and evolve to enable agents’ evolution. The field of psychology already has many rich studies related to this issue(matthews_personality_2003, ; anderson_embodied_2003, ; jr_embodiment_2005, ; solso_cognitive_2005, ). Inspired by the stimulus-response theory in behaviorist psychology(bargh_beyond_2000, ; clark_learning_2018, ), we wish for agents to gain stimulation and feedback through interaction with simulation environments, thereby affecting the development of the Agent’s personality.
Based on the psychology theory discussed(noauthor_experimental_nodate, ; noauthor_classics_nodate, ; vandenbos_apa_2007, ; baldwin_mental_1906, ), we propose a novel agent architecture - Evolving Agents, which encompasses two subsystems: Behavior and Personality, as shown in figure 2. The Behavior system guides the execution of observable, personalized actions. The personality system establishes a cognitive mechanism with individual characteristics capable of continuously updating personality based on feedback from external information. During simulation, the Behavior and Personality systems continuously influence each other. The planning and behavior of the Agent will take its characteristics into careful consideration. Simultaneously, the experiences generated from an agent’s daily actions impact personality through memory and insight. This process of mutual influence forms a feedback loop pattern, facilitating continuous personality evolution.
The behavior system is comprised of two main modules: Plan and Action. Its execution includes personalized planning for the Agent, social engagements and responses, and behavior execution functionalities. The Personality system contains three main modules: Emotion, Cognition, and Character Growth. Its execution involves subjective emotional perception, daily information filtering, memory decay management, comprehensive context-aware idea summarization, and the growth of the Agent’s character. Each module is empowered by large language models like GPT-4 when dealing with its input.
@@ 3.1. Agent Behavior
The Behavior system is responsible for executing planned actions(wang_survey_2024, ) within a simulated environment. Building on the existing architecture of Agent simulation(caron_identifying_2022, ; curry_computer_2023, ; sumers_cognitive_2024, ), which includes planning (Plan) and executing plans (Action)(zhou_least–most_2023, ; song_llm-planner_2023, ; noauthor_230516291_nodate, ; shen_hugginggpt_2023, ), we introduce three innovative elements in the Plan Module: Characteristic Plan, Goal-based Mechanism, and Specific Post-Process. In the Action Module, we propose a new method to describe the Agent’s social relationships and enhance the continuity of conversations between Agents.
@@ 3.1.1. Plan Module
Characteristic Plan: We aim to reflect agents’ personality differences in Evolving Agents’ behaviors. Therefore, in the design of the planning module, we consider not only environmental information as inputs but also integrate agents’ character and memory as references for GPT-4 to generate plans. This mechanism ensures that each Agent can create a schedule aligned with its personality. The specific mechanism by which the Personality system affects the Plan module will be discussed in section 3.2.
@@
Figure 3. Each Agent can create a schedule that aligns with its character. For example, the enthusiastic and open Isabella prefers to leave her dorm and make social plans compared to the shy Benjamin
@@
Goal-based Mechanism: We aim to construct a mechanism that can easily analyze agent behavior and personality variations. Therefore, we referenced relevant perspectives on behavior and motivation from self-determination theory(noauthor_what_nodate, ), considering planning as an expression of behavioral motivation. In our specific design, we defined categories of goals (such as learning, relaxation, exercise, etc.) and required agents to understand and refer to these goals while planning, as shown in Figure 3. This ensures that agents consider the motivations behind each planned behavior, guiding their thought process effectively. Such a goal-based mechanism also aids observers in summarizing user behavior patterns, making it easier to analyze related pattern evolutions. We will discuss this in detail in Chapter 5.
@@
Specific Post-Process: We provide different planning goals with unique post-processes. Taking ”Appointment” as an example, most existing Agent simulation systems use distance-based and probability-based triggering logic for Agent interactions, as is shown in Figure 4. We aim to enhance the proactivity of Agents in interacting with others, fostering more ”conscious” and ”prepared” social activities in the simulated world. In the planning phase, Agents can consider their relationships with other Agents and their current state to decide whether to extend social invitations. The invited Agents respond to the invitation, giving reasons for acceptance or rejection. Agents weigh existing social plans against new ones for overlapping appointments, prioritizing them based on personal development benefits. Following this request-response process, both parties adjust their plans based on the outcome.
@@
For example, Sophia wants to meet Isabella at a cafe at 16:00 to discuss, believing that Isabella’s appreciation for creative expression aligns with her dream of becoming an inspiring and successful writer. Subsequently, Isabella accepts Sophia’s invitation, reasoning that her preference for in-depth conversations over superficial socializing perfectly complements Sophia’s personality, suggesting that their meeting could lead to deeper and more meaningful discussions. Interestingly, when faced with simultaneous social invitations, even though Isabella initially planned to engage with another friend, Benjamin, during this time, she considers the conversation with Sophia potentially more crucial for her personal development and thus accepts this invitation. Please refer to the image for details on the process.
@@ 3.1.2. Action Module
Action Mechanism: When the game time reaches the scheduled plan time, the Agent executes the specific plan. By considering the current plan and personality information, GPT-4-turbo generates a detailed, objective, natural language description of the Agent’s actions regarding the plan. Before executing the plan, the Agent must check if the environmental conditions meet the plan’s objectives, adopting the item interaction condition logic from Generative Agents(park_generative_2023, ). If the target interaction object is occupied, the Agent will replan until conditions are met. Once the environmental object target for executing the plan is secured, the Agent will take over the interaction spot (in social scenarios, reserving adjacent spots for its social peers) and perform the Action.
@@
Continuously Developing Communication: When two Agents come within a certain distance of each other, there’s a probability of triggering a dialogue, depending on their social tendencies and current state (for example, if an Agent’s plan is social, it will actively initiate a conversation). From a social psychology perspective, such interactions allow Agents to gain insights from others with different personalities, serving as a vital external stimulus for their personality development. We aim to gradually increase the depth of conversation between agents, enabling them to generate more valuable insights. To achieve this, we have implemented the dialogue memory. Whenever a conversation occurs, the Agents will review the summary of content previously discussed with the current interlocutor. Then, it determines a ”further” new topic based on the previously discussed topics and discusses this new topic. This dialogue mechanism enhances the inspiration and viewpoint generation. For example, Isabella and Sophia’s discussions deepened from personal development and societal needs to specific methods for maintaining creativity and mental health under social pressure, the procedure of which is shown in Figure 5. This mechanism mirrors real-world communication, where familiarity and discussion depth lead to increasingly inspiring and detailed insights on certain topics.
@@
Figure 5. Agent chooses their communication partner and defines the topic to talk about based on the dialogue memory
@@
Abstract Agent Relationship: Evolving Agents does not use parameters or specific categories to describe agents’ relationship with others(wang_humanoid_2023, ). We found that doing so would excessively predetermine interactions with certain Agents. We enable large language models to select preferred conversation partners based on agents’ dialogue memory. Simultaneously, we require agents to provide reasons for their choices. This approach results in more rational and intriguing social choices, enabling agents to openly choose conversation partners and gain diverse perspectives on their dilemmas. For instance, Isabella, who is balancing external expectations with personal aspirations, might focus on issues like personal dreams and social responsibility with Sophia, a literary writer interested in social insights, or focus on balancing personal pursuits and intimate relationships with Isla, a romantic filmmaker, as is shown in figure 6. This process can help agents explore more possibilities during their Personality Evolution.
@@ 3.2. Agent Personality
In the Agent Behavior System described above, Agents produce a variety of objective behaviors through interactions with their environment and other Agents. From the cognitive psychology perspective, these behaviors can be viewed as stimuli. The Agent Personality system is responsible for processing these stimuli, generating feedback, and thus driving the evolution of the Agent’s personality. The widely accepted definition of personality in psychology encompasses cognitive, emotional, and character traits that influence behaviors(zuckerman_psychobiology_1991, ). We adapt these elements from psychology theory to fit the operational logic of Agent Architecture, ultimately identifying three main modules within the personality system: Emotion, Cognition, and Character Growth.
@@ 3.2.1. Character Structure
Five Dimension-Based Character Structure: In psychology, the character is defined as information that includes habits and tendencies in thought and action(snowden_oxford_nodate, ), which play a crucial role in their behavior. From the system design perspective, a character directly describes an individual’s traits and state.
A structure that can precisely and comprehensively describe characteristics is essential for effectively integrating and transmitting personality information to each module within an agent architecture. We constructed a five-dimension Character Structure based on psychological theories, including Basic Information, Current State, Traits(noauthor_introduction_nodate, ), Conflict(noauthor_apa_nodate, ), and Preference(sen_behaviour_1973, ), as is shown in figure 7. ”Basic Information” encompasses fundamental information about the Agent that does not change, such as gender, age, and professiowetate, we detail the Agent’s initial life and psychologic in the Current Stateal state. Personality Traits describes the Agent based on the Big Five personality dimensions. In conflict, we detail contradictions within the Agent’s personality, facilitating personality evolution around these conflicts. To better describe Agent’s behavior characteristics, in Behavior Preference, we set six-dimension criteria, which are composed of the ultimate goal, long-term goal(derived by an ultimate goal), short-term goal(derived by an ultimate goal), daily routine(based on Agent’s overall personality), hobbies and venue preference. In practice, we initialize specific agents’ personalities using GPT-4-turbo based on this Five Dimension Character structure.
@@
Two Versions of Character Structure: To enhance the efficiency and accuracy of the system development process, we employ two versions of the Character Structure within the system: a full version and a summary version. The full version includes a complete description, while the summary version condenses and summarizes the full version. In specific modules, we differentiate between utilizing the full version and the summary version based on the variance in information weight. For instance, in the Behavior System’s Plan module discussed earlier, we emphasize the information weight of Behavior Preference from the Character Structure. Therefore, we provide a full description of Behavior Preference to GPT-4-turbo but only the summary versions for other dimensions. As for other dimensions, we only provide the summary version. This approach helps the model focus more on the alignment between Plan and Behavior Preference, effectively reducing the generation time for large language models.
@@ 3.2.2. Emotion
Timely Experience and Feeling: Human feelings towards the same objective behavior are not constant. Specific subjective experiences and emotional responses are generated in response to external stimuli under particular circumstances. In specific circumstances, outworld stimuli affect us to generate emotional behavior and subjective feelings. As a result, our cognitive process, such as attention, thinking patterns and decisions, is influenced, leading us to differentiated perceptions of external information(solso_cognitive_2005, ). Therefore, it’s crucial to build a system that can not only comprehensively consider the current state and perform emotional perception on ongoing behaviors, but also ensure that the later cognition processes receive information aligned with the character’s current emotional state. At the implementation level, we categorize the Agent’s emotional state into seven categories. The Emotion module considers the current personality conflicts and other personality information to update emotions based on the behavior being executed. In detail, we request that GPT4 choose the current emotion class with a first-person perspective description of inner feelings. This module allows Agents to generate subjective descriptions that vary and align with their experiences, such as Sophia feeling excited about writing her novel due to anticipation of creating her literary work or feeling anxious and self-doubtful after hearing the technological innovation in the writing process told by Benjamin, as is shown in Figure 8.
@@
Emotion-Based Plan Evaluation: Previous research has noted that humans don’t rigidly follow the plans they set for themselves(minsky_society_1988, ). In the design of Humanoid Agent architecture, researchers proposed a dual-system structure of emotion and behavior to allow Agents to adapt their plans based on internal feelings constantly.
The Evolving Agents’ Emotion module supports continuous evaluation and adjustment in the Behavior system’s Plan module. We simplified the adjustment mechanism of Emotion on the Plan, using a more systematic decision-making mechanism to trigger plan adjustments. Emotions are categorized on a scale from negative to positive, mapped within a 0-7 interval. Significant emotional shifts during an action (e.g., from excitement to calm, crossing a 3-interval span) indicate substantial emotional fluctuations, resulting in which agents should reevaluate the forthcoming plans. This straightforward strategy, as shown in Figure 9, demonstrates believable and dynamic planning capabilities. For example, after a conversation with Sophia, Isabella’s mood shifted from excitement to fear (due to the discussion regarding social responsibilities caused pressure), leading her to adjust her original plan and go to a square for relaxation to calm her emotions.
Figure 9. Agent evaluates and modifies existing plans based on the emotional fluctuations